OpenStudy (konradzuse):

What is the required limit of lim ->0+ ...(sin(x))^x ?

OpenStudy (anonymous):

uh, what? \[\lim_{x \to 0}+\dots \sin^x(x)\]?

OpenStudy (konradzuse):

Yeah it's asking "what is the required limit" I have no idea what they are asking...

OpenStudy (anonymous):

I have no idea what the expression even means. It isn't written using (standard) legible notation.

OpenStudy (konradzuse):

Without the .... obviously... I was doing that so it didn't look like 0+ (sin(x))^x

OpenStudy (konradzuse):

\[\lim_{x \rightarrow 0+} (\sin(x))^x\]

OpenStudy (anonymous):

That's more clear. That's actually solvable. I think it's asking for the right-hand limit.

OpenStudy (turingtest):

I agree^ LaTeX note: \lim_{x\to0^+} \[ \lim_{x\to0^+}\]that makes it more clear :)

OpenStudy (turingtest):

is there a better way to do this than doing the exponent trick?

OpenStudy (anonymous):

I'm confused as to why we can't evaluate this simply. \[\lim_{x \to 0^{+}}\sin^x(x)=sin^0(0)=(0)^0=1?\]

OpenStudy (turingtest):

0^0 is undefined

OpenStudy (konradzuse):

I got that, but I'm not sure what they mean by the "limit." and what we are trying to solve for...

OpenStudy (anonymous):

@TuringTest, no. It is conventionally defined. Most authors take \(0^0=1\).

OpenStudy (turingtest):

Except for by some mathematicians like Euler who took it upon themselves to define 0^0=1 I beg to differ. The truth is that the issue is still very debated. But don't take my word for it, let me find a good link I saw once, one sec...

OpenStudy (anonymous):

I have read about it. It's generally a convention/definition. See http://www.askamathematician.com/2010/12/q-what-does-00-zero-raised-to-the-zeroth-power-equal-why-do-mathematicians-and-high-school-teachers-disagree/

OpenStudy (anonymous):

But we are detracting from the question here. Is that the actual answer?

OpenStudy (turingtest):

I dunno, let's just exponentiate

OpenStudy (konradzuse):

yes I just ented 1 and i was correct.

OpenStudy (konradzuse):

I got that before, but I dind't understand exactly what they were asking... Solve for the limit wuld havebeen simpler.. I thought they were asking to change the imit to something else which made no sense.. Thanks Limitless...

OpenStudy (anonymous):

You're welcome, @KonradZuse! :) Also, this is freaky: http://www.wolframalpha.com/input/?i=lim+as+x+to+0%5E%2B+of+sin%5Ex%28x%29

OpenStudy (konradzuse):

http://www.wolframalpha.com/input/?i=limit+x+-%3E+0%2B++of+%28sin%28x%29%29%5Ex This is what I had :P

OpenStudy (konradzuse):

Pool time everyone, ttyl.

OpenStudy (anonymous):

@KonradZuse, ah! WolframAlpha failed to interpret it. Nice catch.